# AI-enhanced Collective Intelligence: The State of the Art and Prospects

### v2

@@
Abstract.
@@
The current societal challenges exceed the capacity of human individual or collective effort alone. As AI evolves, its role within human collectives is poised to vary from an assistive tool to a participatory member. Humans and AI possess complementary capabilities that, when synergized, can achieve a level of collective intelligence that surpasses the collective capabilities of either humans or AI in isolation. However, the interactions in human-AI systems are inherently complex, involving intricate processes and interdependencies. This review incorporates perspectives from network science to conceptualize a multilayer representation of human-AI collective intelligence, comprising a cognition layer, a physical layer, and an information layer. Within this multilayer network, humans and AI agents exhibit varying characteristics; humans differ in diversity from surface-level to deep-level attributes, while AI agents range in degrees of functionality and anthropomorphism. The interplay among these agents shapes the overall structure and dynamics of the system. We explore how agents’ diversity and interactions influence the system’s collective intelligence. Furthermore, we present an analysis of real-world instances of AI-enhanced collective intelligence. We conclude by addressing the potential challenges in AI-enhanced collective intelligence and offer perspectives on future developments in this field.
@@
AI, Collective Intelligence, Hybrid Intelligence, Multi-agent systems, Human-machine networks
1. Introduction
@@
Our societies form, sustain, and function because of our intelligence. In the animal kingdom, the more intelligent the animals are, the more social they are with each other. Dunbar’s ”social brain theory” (Dunbar, 1998) even suggests that significantly more superior human intelligence compared with other primates is the result of our need to be able to manage and maintain our social lives; it is humans’ outstanding sociality that forced them to become more intelligent. The bottom line is that in human societies, collectives are intelligent, and intelligence is collective. As a concept, collective Intelligence (CI) refers to the emergent outcome of the collective efforts of many individuals. The superiority of collective intelligence to the intelligence of any of the individuals who contributed to it has been demonstrated scientifically and practically in numerous instances (Bonabeau, 2009).
@@
In every new chapter in the history of Information and Communication Technologies (ICT), collective intelligence has been elevated to a higher level through more readily and cheaply available platforms for collaboration and exchanging ideas. The Internet, in general, and in particular, the World Wide Web, was designed to foster large-scale collaboration. Tim Berners-Lee’s initial intention in developing HTTP, which later led to the development of the WWW, was to facilitate collaboration between CERN researchers. Some thirty years later, the Web facilitates the largest collaborative projects in human history, such as Wikipedia (Yasseri and Kertész, 2013), citizen science projects (Ibrahim et al., 2021), collaborative software development, and collaborative tagging projects (Yasseri and Menczer, 2023).
@@ 1.1. Collective intelligence
Intelligent agents (humans, animals, and intelligent artifacts) often interact with others to tackle complex problems. They combine their knowledge and personal information with information acquired through interaction and social information, resulting in superior collective intelligence. The concept of collective intelligence has been redefined over the years, primarily due to different research streams that have employed the term in qualitatively different ways (Woolley and Gupta, 2023).
One stream of research uses collective intelligence to refer to the outcome of collaboration produced by amalgamation of the input from a large crowd, exemplified in scenarios such as online contests or crowdsourced science (Aristeidou and Herodotou, 2020; Kittur et al., 2009; Malone, 2018b). Here, the concept of collective intelligence is often mixed up with the wisdom of the crowd. The wisdom of the crowd refers to judgments achieved by large groups and crowds consisting of distributed individuals through the pooling of personal information, often via simple averaging (Surowiecki, 2004). It relies on the statistical aggregation of independent judgments to collectively arrive at a decision or prediction that is often better than the answer given by any single member of the group. In contrast, collective intelligence is more related to achievements by smaller collaborating groups and pairs, traditionally studied within the broader topic of group performance and decision-making (Kerr and Tindale, 2004; Malone et al., 2010). A more theoretical distinction is based on the fact that collective intelligence emerges from the synergy of interactions between intelligent agents (Grasso and Convertino, 2012) collaborating towards a common goal through interconnected efforts.
@@
Another stream of research defines collective intelligence as an ability that ”can be designed to solve a wide range of problems over time in the face of complex and changing environmental conditions” (Woolley and Gupta, 2023). Converging evidence suggests the presence of a general collective intelligence factor (”c-factor”) that serves as a predictor for a group’s performance across a wide range of tasks  (Woolley et al., 2010, 2015), similar to the individual’s general intelligence (g-factor) but extended to groups. Research has found that this ”c-factor” is not strongly correlated with the average or maximum individual intelligence of the group members, suggesting that collective intelligence is more than just the sum of the individual intelligence present in the group (Woolley et al., 2010). In this sense, this definition is virtually compatible with the one above.
@@
A group’s collective intelligence is fundamentally influenced by its composition and interactions. Past work indicates a positive correlation between the ”c-factor” and factors such as the average social sensitivity  (Woolley et al., 2010), the equitable distribution of conversational participation  (Woolley et al., 2010), and the percentage of female members in the group  (Woolley et al., 2010). Groups with diverse perspectives and skill sets are likelier to cultivate innovative solutions that might not emerge in a more homogenous group. Effective interaction processes, encompassing coordination, cooperation (Malone and Bernstein, 2022), and communication patterns (Engel et al., 2015) are crucial. Prior studies have identified three distinct socio-cognitive transactive systems  (Woolley et al., 2023a) responsible for managing collective memory (Yasseri et al., 2022), collective attention (Gildersleve et al., 2023), and collective reasoning (Wyss and Beste, 2017), all of which are essential for the emergence of collective intelligence.
1.2. Artificial intelligence
@@
The definition of artificial intelligence has undergone numerous evolutions in the past semi-century. Artificial intelligence can be understood as the simulation of human intelligence processes by machines (Konar, 2018), especially computer systems. Previous research considers the classification of AI into two distinct categories (Hassani et al., 2020). The first classification categorizes AI based on the level of human-like cognitive abilities they exhibit, such as the ability to ”think” and ”feel.” There are four primary AI types (Burns et al., 2021): reactive AI, limited memory AI, theory of mind AI, and self-aware AI. Among them, the limited-memory AI is the most common at the current time, ranging from virtual assistants to chatbots to self-driving vehicles. Such AI can learn from historical data to recognize patterns, generate new knowledge and understanding, and inform subsequent decisions.
@@
The second classification is a technology-oriented approach that categorizes AI into Artificial Narrow Intelligence (ANI), Artificial General Intelligence (AGI), and Artificial Superintelligence (ASI). ANI involves creating computer systems capable of performing specific tasks in a manner similar to human intelligence (Rawat et al., 2023), but often surpass humans in efficiency and accuracy (Kuusi and Heinonen, 2022). ANI machines have a narrow range of capabilities and can represent all existing AI. AGI refers to machines that exhibit the human ability to learn, perceive, and understand a wide range of intellectual tasks (Goertzel, 2014; Review, 2023). Finally, the primary goal of ASI is to develop a machine that exhibits cognitive abilities higher than those of humans (Hassani et al., 2020).
@@
Existing AI can perform tasks including but not limited to complex calculations, language translation, facial recognition, and financial market prediction. Recent development in generative AI, such as OpenAI’s text-to-video model Sora (OpenAI, 2024), exemplifies AI’s potential in creative industries. As AI technology evolves, it is poised to bring groundbreaking advancements across various fields. Beyond content creation, as AI deepens the understanding of the physical world and develops its simulation ability at the perceptual and cognitive level, it is possible to develop more super-intelligent tools in various other fields.
@@
AI’s ability to process vast amounts of data on a scale is far beyond human capabilities. However, due to its limitations, many real-world applications cannot yet be solved solely by AI. AI currently lacks the deep conceptual and emotional understanding that humans possess about objects and experiences (De Cremer and Narayanan, 2023). AI cannot interpret human language’s nuances, including contextual, symbolic, and cultural meanings (Nath and Sahu, 2020; Mitchell, 2021). AI cannot be an ethical decision-maker due to a lack of human attributes of intentionality, care, and responsibility (De Cremer and Narayanan, 2023). Therefore, combining human insight with AI’s analytical power is crucial for addressing complex real-world challenges, leveraging the strengths of both to compensate for each other’s weaknesses.
1.3. Human-AI hybrid collective intelligence
@@
Even though the field of collective intelligence initially focused solely on groups of people, in recent years, it has gradually expanded to include AI as group members in a new framework referred to as ”hybrid intelligence” (Dellermann et al., 2019). Proponents of the hybrid intelligence perspective stress that humans and AI can connect in ways that allow them to collectively act more intelligently and achieve goals that are unreachable by any of the individual entities alone (Akata et al., 2020). Researchers investigating this hybrid collective intelligence explore ”how people and computers can be connected so that collectively they act more intelligently than any person, group, or computer has ever done before” (Malone, 2018a).
@@
Acknowledging the complementary capabilities of humans and AI as discussed in the previous section, researchers identify the need for developing socio-technological ensembles of humans and intelligent machines that possess the ability to collectively achieve superior results and continuously improve by learning from each other (Dellermann et al., 2021). Previous research provides evidence that teaming humans with AI has the potential to achieve complementary team performance (CTP), a level of performance that AI or humans cannot reach in isolation (McNeese et al., 2021b). For instance, a study demonstrates that humans can use contextual information to adjust the AI’s decision (Hemmer et al., 2022). Research on mixed teams composed of both humans and AI shows hybrid teams could achieve higher performance metrics, such as team situational awareness and team score, compared to all-human teams (McNeese et al., 2021b).
1.4. Empirical evidence and theoretical gaps
@@
Technological advances have constantly disrupted how we produce, exchange, collect, and analyze information and consequently make decisions, individually or collectively. Automation through smart devices based on machine learning, knowledge graph-based machine reasoning, and walking and talking household devices leveraging natural language processing have already reshaped our personal, professional, and social lives. Collective intelligence cannot be an exception. Inevitably, our collective decision-making processes have been, and will be, disrupted by AI.
@@
Machine learning and automation can increase the efficiency and scalability of collective intelligence in Citizen Science projects, where citizens volunteer to help scientists tag and classify their large-scale datasets (Jackson et al., 2018). Members of a crowd with different interests and areas of competence can be matched to other tasks by recommendation systems (Yuen et al., 2011). Generative algorithms can be used to extrapolate human solutions and generate new ideas (Scirea et al., 2015). Clustering algorithms can be used to reduce the solution space of a complex task as humans are exploring possible solutions (Hocking et al., 2018). Machines can unify similar solutions to mitigate statistical noise (Zhang et al., 2015). Matching algorithms can be used to match individuals and build efficient groups (Karger et al., 2011).
@@
This review focuses on the integration of AI to bolster the collective intelligence of human groups, addressing a notable theoretical gap in understanding performance enhancements, particularly in hybrid human-AI configurations. With AI integration becoming pervasive across various sectors, exploring how this collaboration can unlock optimal capabilities is imperative. This exploration is essential not only for boosting productivity and fostering innovation but also for ensuring that AI complements and enhances human skills rather than replacing them outright. The industry’s pivotal role in driving current AI-CI research underscores the absence of a comprehensive theoretical framework (Berditchevskaia et al., 2022).
@@
In subsequent sections, we introduce a dedicated framework to deepen the comprehension of AI-enhanced collective intelligence systems. We then delve into analyzing applications and discussing the implications of such fusion. Finally, we conclude by addressing existing challenges in designing AI-CI systems and speculate on the future trajectory of the field.
2. AI-enhanced collective intelligence framework
2.1. Multilayer representation of the human-AI complex system
@@
In advancing the field of human-AI collective intelligence, prior research shows the necessity of formulating theories encompassing collective intelligence, combining both humans and AI (Eide et al., 2016). Developing such theoretical frameworks requires an in-depth comprehension and interpretation of human-AI systems, which are characterized by their high complexity and interrelated processes. Network science offers tools that enable us to understand the complexity of social systems (Pedreschi et al., 2023) where a multitude of interacting elements give rise to the collective behaviors of the whole system (Bianconi et al., 2023). To enhance our understanding of collective intelligence in the human-AI system, and explore how ”the whole is more than the sum of the individual parts” (Aristotle and Aristotle, 1933; Anderson, 1972), we integrate complexity science and network science approaches, and propose a multilayer representation of the complex system involving human and AI agents.
@@
Figure 1. A multilayer representation to untangle the processes in the complex system of human and AI agents, with three layers influencing each other: the cognition layer, the physical layer, and the information layer. External factors and the changing environment can influence the collective intelligence of the whole system.
@@
As shown in Fig. 1, a real-world collective intelligence system can be mapped into a multilayer network setting where three layers influence each other: cognition, physical, and information. The interacting agents are represented as nodes, and the relations are represented as links connecting the nodes in the network. The construction of such networks and the meaning of the links are context-dependent. A node can also represent a group of agents instead of an individual agent in situations where the group behaves or is treated as a single entity within the network. Across layers, the nodes are the same, but the interactions are different.
@@
The cognition layer is only indirectly perceivable, which contains mental processes. Intelligent processes are involved during problem-solving, such as sense-making, remembrance, creation, coordination, and decision-making, often happening in the cognition layer. The links are hard to determine, but these processes exist and are fundamental to the emergence of collective intelligence in the whole system. In opposition to this indirectly perceivable layer, there is the physical layer where humans and AI have tangible physical interactions. Information interaction in the information layer refers to exchanging information between agents through various communication channels. Even though our current framework revolves around pairwise interactions between nodes, recent advancements in the study of higher-order networks, delving beyond pairwise interactions (Battiston et al., 2020), can significantly enrich our understanding of the collective emergence of intelligence.
@@
In addition to internal processes, it is essential to acknowledge the influence of the environment on the system (Janssens et al., 2022). The human-AI complex system functions in a potentially dynamic environment. Previous research suggests that collective intelligent behavior depends on the environment (Janssens et al., 2022), as certain behaviors or strategies that work well in one setting may not be as effective or relevant in another (Kämmer et al., 2014).
@@
Human-AI hybrid systems can be viewed as complex adaptive systems, continually evolving and adapting through interactions within dynamic environments (Goldstein, 2011). The concept of emergence in complex systems can be used to describe the phenomenon where new properties emerge at the collective level, which is not present at the level of individual components (Goldstein, 2011; Meehl and Sellars, 1956; Girvin and Yang, 2019). Here, the collective intelligence of the whole system is such an emergent property superior to the micro-level individual intelligence. This property, which includes outcomes, abilities, characteristics, and behaviors, is aligned with and encompasses the major existing definitions of collective intelligence. It emerges through complex nonlinear relationships between the agents and is likely a result of both bottom-up and top-down processes (Woolley et al., 2015). The former encompasses the aggregation of group-member characteristics that foster collaboration. The latter includes structures, norms, and routines that govern collective behavior, thereby influencing the efficacy of coordination and collaboration (Woolley et al., 2015). Employing a multilayer representation of this complex system in a changing environment can facilitate a deeper understanding of the interactions and relationships between human and AI agents, untangling the intricate processes involved in the emergence and maintenance of collective intelligence.
2.2. The goal and the task
@@
Task types can vary significantly, from generative activities such as brainstorming, which requires creative and divergent thinking, to analytical tasks such as solving Sudoku, which demands logical reasoning and pattern recognition (Delahaye, 2006). The diversity of the tasks requires different abilities, including memory utilization, creative imagination, sense-making, and critical analysis. In the following subsections, we delve into details of group-member diversity’s influence, the structure and dynamics of interaction networks, and AI’s role in enhancing group performance.
2.3. Diversity
The relationship between diversity and group collective intelligence is highly complex due to the richness of the notion of diversity and its effects under different conditions. A group can exhibit surface-level and deep-level diversities (Phillips and Loyd, 2006). The surface-level diversity comes from readily observable social categories, including gender, age, and ethnicity (Wegge and Schmidt, 2009). The deep-level diversity refers to differences in psychological characteristics (Harrison et al., 2002), and can include personality, cognitive thinking styles, and values.
Diversity has shown benefits for work teams in terms of more creativity, improved decision-making, and access to a broader pool of talent  (Bagshaw, 2004; Van der Vegt and Janssen, 2003). However, diversity can also have adverse effects, such as emotional conflict (Pelled et al., 1999; Yasseri et al., 2012b), stress (DeFrank and Ivancevich, 1998), poor work relationships, and poor overall performance (Van Knippenberg et al., 2004). There is a lack of consensus on which aspects of group diversity are likely to result in positive outcomes and which aspects could produce negative outcomes (Van Knippenberg et al., 2004; Jansen and Searle, 2021). The effects of diversity on group performance depend on the complexity of the task and various moderators, such as the density of the interaction network (Baumann et al., 2024).
2.3.1. Surface-level diversity

### v3

@@
AI-enhanced collective intelligence
Hao Cui School of Sociology, University College Dublin, Belfield, Dublin 4, Dublin, Ireland Geary Institute for Public Policy, University College Dublin, Belfield, Dublin 4, Dublin, Ireland Taha Yasseri School of Sociology, University College Dublin, Belfield, Dublin 4, Dublin, Ireland Geary Institute for Public Policy, University College Dublin, Belfield, Dublin 4, Dublin, Ireland
Summary
@@
The current societal challenges exceed the capacity of human individual or collective effort alone. As AI evolves, its role within human collectives is poised to vary from an assistive tool to a participatory member. Humans and AI possess complementary capabilities that, when synergized, can achieve a level of collective intelligence that surpasses the collective capabilities of either humans or AI in isolation. However, the interactions in human-AI systems are inherently complex, involving intricate processes and interdependencies. This narrative review incorporates perspectives from complex network science to conceptualize a multilayer representation of human-AI collective intelligence, comprising cognition, physical, and information layers. Within this multilayer network, humans and AI agents exhibit varying characteristics; humans differ in diversity from surface-level to deep-level attributes, while AI agents range in degrees of functionality and anthropomorphism. The interplay among these agents shapes the overall structure and dynamics of the system. We explore how agents’ diversity and interactions influence the system’s collective intelligence. Furthermore, we present an analysis of real-world instances of AI-enhanced collective intelligence. We conclude by addressing the potential challenges in AI-enhanced collective intelligence and offer perspectives on future developments in this field.
Keywords
@@
AI, Collective Intelligence, Hybrid Intelligence, Multi-agent systems, Human-machine networks, Human-machine Intelligence
Introduction
@@
Our societies form, sustain, and function because of our intelligence. In the animal kingdom, the more intelligent the animals are, the more social they are. Dunbar’s “social brain theory” 1 even suggests that significantly more superior human intelligence compared with other primates is the result of our need to be able to manage and maintain our social lives; it is humans’ outstanding sociality that forced them to become more intelligent. The bottom line is that in human societies, collectives are intelligent, and intelligence is collective. As a concept, Collective Intelligence (CI) refers to the emergent outcome of the collective efforts of many individuals. The superiority of collective intelligence to the intelligence of any of the individuals who contributed to it has been demonstrated scientifically and practically in numerous instances 2.
@@
In every new chapter in the history of Information and Communication Technologies (ICT), collective intelligence has been elevated to a higher level through more readily and cheaply available platforms for collaboration and exchanging ideas. The Internet, especially the World Wide Web, was designed to foster large-scale collaboration. Tim Berners-Lee’s initial intention in developing HTTP, which later led to the development of the WWW, was to facilitate collaboration between CERN researchers. Some thirty years later, the Web facilitates the largest collaborative projects in human history, such as Wikipedia 3, citizen science projects 4, collaborative software development, and collaborative tagging projects 5, all exemplifying technology-enhanced collective intelligence
@@
Intelligent agents (humans, animals, and intelligent artifacts) often interact with others to tackle complex problems. They combine their knowledge and personal information with information acquired through interaction and social information, resulting in superior collective intelligence. The concept of collective intelligence has been redefined over the years, primarily due to different research streams that have employed the term in qualitatively different ways 6.
Aspect Wisdom of Crowds Collective Intelligence
Interaction Minimal High, involving significant collaboration and coordination.
Dependence Independent Interdependent and adaptable processes to changing conditions.
Mechanism Aggregation Emergent and adaptable integration from coordinated efforts.
Nature Static, immediate collective judgment Dynamic and adaptive, evolving to meet complex needs.
Table 1: Comparison between Wisdom of Crowds and Collective Intelligence
One stream of research uses collective intelligence to refer to the outcome of collaboration produced by amalgamation of the input from a large crowd, exemplified in scenarios such as online contests or crowdsourced science 7, 8, 9. Here, the concept of collective intelligence is often mixed up with the wisdom of crowds (WoC). Both rely on the idea that collective input from a diverse group typically yields better outcomes than any individual member’s. The wisdom of crowds involves aggregating independent judgments from a large, diverse group to arrive at collective decisions or predictions, often through statistical averaging 10. In contrast, collective intelligence pertains to the achievements of collaborating groups, whose sizes can vary. It emerges from the synergy of interactions and mutual feedback among intelligent agents 11 working towards a common goal through interconnected efforts. However, social information and interconnection have also been studied in the context of WoC 12, 13. Despite the conflict in the literature on the boundaries and overlaps between the two, almost all the AI enhancements we discuss in this paper in the context of CI can also be implemented in WoC scenarios 14. Nevertheless, for the sake of simplicity, we only use the term collective intelligence for the rest of the paper. Table. 1 compares key aspects of the two concepts.
@@
Another stream of research defines collective intelligence as an ability that “can be designed to solve a wide range of problems over time in the face of complex and changing environmental conditions” 6. Converging evidence suggests the presence of a general collective intelligence factor (“c-factor”) that serves as a predictor for a group’s performance across a wide range of tasks  15, 16, similar to the individual’s general intelligence (g-factor) but extended to groups. Research has found that this “c-factor” is not strongly correlated with the average or maximum individual intelligence of the group members, suggesting that collective intelligence is more than just the sum of the individual intelligence present in the group 15. In this sense, this definition is virtually compatible with the one above.
@@
A group’s collective intelligence is fundamentally influenced by its composition and interactions. Past work indicates a positive correlation between the “c-factor” and factors such as the average social sensitivity  15, the equitable distribution of conversational participation  15, and the percentage of female members in the group  15. Groups with diverse perspectives and skill sets are likelier to cultivate innovative solutions that might not emerge in a more homogenous group. Effective interaction processes, encompassing coordination, cooperation 17, and communication patterns 18 are crucial. Prior studies have identified three distinct socio-cognitive transactive systems  19 responsible for managing collective memory 20, collective attention 21, and collective reasoning 22, all of which are essential for the emergence of collective intelligence.
Artificial intelligence
@@
The definition of artificial intelligence has undergone numerous evolutions in the past semi-century. Artificial intelligence is the simulation of human intelligence processes by machines 23, especially computer systems. Previous research considers the classification of AI in distinct ways 24. The first classification categorizes AI based on its human-like cognitive abilities, such as “think” and “feel.” In another categorization, there are four primary AI types 25: reactive AI, limited memory AI, theory of mind AI, and self-aware AI. Limited-memory AI is the most common at the current time, ranging from virtual assistants to chatbots to self-driving vehicles. Such AI can learn from historical data to recognize patterns, generate new knowledge and understanding, and inform subsequent decisions.
@@
Yet, another classification is a technology-oriented approach that categorizes AI into Artificial Narrow Intelligence (ANI), Artificial General Intelligence (AGI), and Artificial Superintelligence (ASI). ANI involves creating computer systems capable of performing specific tasks like human intelligence 26, but often surpass humans in efficiency and accuracy 27. ANI machines have a narrow range of capabilities and can represent all existing AI. AGI refers to machines that exhibit the human ability to learn, perceive, and understand a wide range of intellectual tasks 28, 29. Finally, ASI’s primary goal is to develop a machine with cognitive abilities higher than those of humans 24.
@@
Existing AI can perform tasks including but not limited to complex calculations, language translation, facial recognition, and financial market prediction. Recent developments in generative AI, such as OpenAI’s text-to-video model Sora 30, exemplify AI’s potential in creative industries. As AI technology evolves, it is poised to bring groundbreaking advancements across various fields. Beyond content creation, as AI deepens the understanding of the physical world and develops its simulation ability at the perceptual and cognitive level, it is possible to develop more super-intelligent tools in various fields.
@@
Current AI systems can process vast amounts of data on a scale far beyond human capabilities. However, many real-world challenges cannot yet be solved solely by AI. Currently, AI lacks the deep conceptual and emotional understanding humans possess about objects and experiences 31. AI cannot interpret human language’s nuances, including contextual, symbolic, and cultural meanings 32, 33. AI cannot be an ethical decision-maker because it lacks the human attributes of intentionality, care, and responsibility 31. Therefore, combining human insight with AI’s analytical power is crucial for addressing complex real-world challenges, leveraging the strengths to compensate for each other’s weaknesses.
Human-AI hybrid collective intelligence
@@
Even though the field of collective intelligence initially focused solely on groups of people, in recent years, it has gradually expanded to include AI as group members in a new framework referred to as “hybrid intelligence” 34. Proponents of the hybrid intelligence perspective stress that humans and AI can connect in ways that allow them to collectively act more intelligently and achieve unreachable goals by any individual entities alone 35. Researchers investigating this hybrid collective intelligence explore “how people and computers can be connected so that collectively they act more intelligently than any person, group, or computer has ever done before” 36.
@@
Acknowledging the complementary capabilities of humans and AI as discussed in the previous section, researchers identify the need for developing socio-technological ensembles of humans and intelligent machines that possess the ability to collectively achieve superior results and continuously improve by learning from each other 37. Previous research provides evidence that teaming humans with AI has the potential to achieve complementary team performance (CTP), a level of performance that AI or humans cannot reach in isolation 38. For instance, a study demonstrates that humans can use contextual information to adjust the AI’s decision 39. Research on mixed teams composed of humans and AI shows hybrid teams could achieve higher performance metrics, such as team situational awareness and score, than all-human teams 38.
Empirical evidence, theoretical gaps, and a new framework
@@
Technological advances have constantly disrupted how we produce, exchange, collect, and analyze information and consequently make decisions, individually or collectively. Automation through smart devices based on machine learning, knowledge graph-based machine reasoning, and walking and talking household devices leveraging natural language processing, have already reshaped our personal, professional, and social lives. Collective intelligence cannot be an exception. Inevitably, our collective decision-making processes have been, and will be, disrupted by AI.
@@
For example, machine learning and automation can increase the efficiency and scalability of collective intelligence in Citizen Science projects, where citizens volunteer to help scientists tag and classify their large-scale datasets 40. Members of a crowd with different interests and areas of competence can be matched to other tasks by recommendation systems 41. Generative algorithms can extrapolate human solutions and generate new ideas 42. Clustering algorithms can reduce a complex task’s solution space as humans explore possible solutions 43. Machines can unify similar solutions to mitigate statistical noise 44. Matching algorithms can match individuals and build efficient groups 45.
In this review of AI-enhanced collective intelligence, we adopted a narrative review methodology guided by a complexity theory and network science framework. We began with a conceptual narrative and systematically identified relevant literature for each section. Therefore, we do not claim we have exhausted the vast and fast-growing literature. However, our approach integrates diverse theoretical perspectives and empirical studies, using these theories to explain collective intelligence and its enhancement by AI. This approach bridges interdisciplinary insights, offering a holistic understanding of AI-CI.
@@
This review focuses on integrating AI to bolster the collective intelligence of human groups, addressing a notable theoretical gap in understanding performance enhancements, particularly in hybrid human-AI configurations. With AI integration becoming pervasive across various sectors, exploring how this collaboration can unlock optimal capabilities is imperative. This exploration is essential for boosting productivity, fostering innovation, and ensuring that AI complements and enhances human skills rather than replacing them outright. The industry’s pivotal role in driving current AI-CI research underscores the absence of a comprehensive theoretical framework 46.
@@
In subsequent sections, we introduce a framework to deepen our comprehension of AI-enhanced collective intelligence systems. We then elaborate on the applications and discuss the implications of such fusion. Finally, we conclude by addressing existing challenges in designing AI-CI systems and speculate on the field’s future trajectory.
AI-enhanced collective intelligence framework
Multilayer representation of the collective intelligence system
@@
In advancing the field of human-AI collective intelligence, prior research shows the necessity of formulating theories encompassing collective intelligence, combining both humans and AI 47. Developing such theoretical frameworks requires an in-depth comprehension and interpretation of human-AI systems, characterized by their high complexity and interrelated processes. Network science offers tools that enable us to understand the complexity of social systems 48 where a multitude of interacting elements give rise to the collective behavior of the whole system 49. To enhance our understanding of collective intelligence in the human-AI system and explore how “the whole is more than the sum of its parts” 50, 51, we integrate complexity science and network science approaches and propose a multilayer representation of the complex system involving human and AI agents.
@@
Figure 1: A multilayer representation to untangle the processes in the complex system of human and AI agents, with three layers influencing each other: the cognition, physical, and information layers. External factors and the changing environment can influence the emergent collective intelligence of the whole system.
@@
Complex system thinking has been used to understand diverse biological, physical, and social domains 52, 53, 54. Inspired by this approach, a real-world collective intelligence system can be mapped into a multilayer network with three interconnected layers: cognition, physical, and information. In this network, nodes represent interacting agents, and links represent relationships between them. The construction of such networks and the meaning of the links are context-dependent. A node can also represent a group of agents instead of an individual agent in situations where the group behaves or is treated as a single entity within the network. Figure 1 illustrates a multiplex network 55, a special type of multilayer network where nodes remain the same but links differ across layers.
@@
The cognition layer, which contains mental processes, is only indirectly perceivable. Intelligent processes are involved during problem-solving, such as sense-making, remembrance, creation, coordination, and decision-making, often happening in the cognition layer. The links are hard to determine, but these processes exist and are fundamental to the emergence of collective intelligence in the whole system. In opposition to this indirectly perceivable layer is the physical layer, where humans and AI have tangible physical interactions. Information interaction in the information layer refers to exchanging information between agents through various communication channels. Besides intra-layer interactions, interactions in one layer can lead to information transfer and trigger interactions in other layers, represented as inter-layer links in the network 56. This interdependence and cross-layer influence are essential features of multilayer networks. Moreover, Even though our current framework revolves around pairwise interactions between nodes, recent advancements in the study of higher-order networks, going beyond pairwise interactions 57, may enrich our understanding of the collective emergence of intelligence.
@@
In addition to internal processes, it is essential to acknowledge the influence of the environment on the system 58. The human-AI complex system functions in a potentially dynamic environment. Previous research suggests that collective intelligent behavior depends on the environment 58, as certain behaviors or strategies that work well in one setting may not be as effective or relevant in another 59.
Components Features
Layers Cognition layer Information layer Physical layer
Nodes Human: surface-level, deep-level diversity AI agents: diversity in functionality and anthropomorphism
Links Links can have directions. Intra-layer links: interactions within layers Inter-layer links: interactions between layers
Structure Size, centrality, density, hierarchy, community …
Dynamics Communication patterns Cognitive processes etc
Table 2: Summary and breakdown of the multilayer network of the human-AI complex system.
@@
Human-AI hybrid systems can be viewed as complex adaptive systems, continually evolving and adapting through interactions within dynamic environments 60. The concept of emergence in complex systems can be used to describe the phenomenon where new properties emerge at the collective level, which is not present at the level of individual components 61, 60, 62, 63. Here, the collective intelligence of the whole system can be seen as an emergent property superior to the micro-level individual intelligence. This property, which includes outcomes, abilities, characteristics, and behaviors, is aligned with and encompasses the major existing definitions of collective intelligence. It emerges through complex nonlinear relationships between the agents and is likely a result of bottom-up and top-down processes 16. The former encompasses the aggregation of group-member characteristics that foster collaboration. The latter includes structures, norms, and routines that govern collective behavior, thereby influencing the efficacy of coordination and collaboration 16. The modeling of the emergence of collective intelligence has leveraged analogies from a diverse array of fields 64, ranging from statistical physics 65 to neuroscience 66, both within and beyond.
Employing a multilayer representation of this complex system in a changing environment can facilitate a deeper understanding of the interactions and relationships between humans and AI agents, untangling the intricate processes involved in emerging and maintaining collective intelligence. Given the interdisciplinary nature of this approach and the variety of terminologies used, key terms and main concepts are listed in supplemental information Table S1 for more clarity. Table 2 provides a summary and breakdown of the components and features within the proposed multilayer network. The immediate benefit of this framework is that we can learn from the vast literature on multi-layer networks and the studies on their robustness67, adaptivity68, scalability69, resielence70, and interoperability71. We will discuss this further in the following sections.
The goal and the task
@@
Task types can vary significantly, from generative activities such as brainstorming, which requires creative and divergent thinking, to analytical tasks such as solving Sudoku, which demands logical reasoning and pattern recognition 72. The diversity of the tasks requires different abilities, including memory utilization, creative imagination, sense-making, and critical analysis. In the following subsections, we focus on the details of group-member diversity’s influence, the structure and dynamics of interaction networks, and AI’s role in enhancing group performance.
Group diversity
The relationship between group diversity and collective intelligence is highly complex due to diversity’s complexity and the range of its effects under different conditions. A group can exhibit surface-level diversities 73. The surface-level diversity comes from readily observable social categories, including gender, age, and ethnicity 74. The deep-level diversity refers to differences in psychological characteristics 75, including personality, cognitive thinking styles, and values.
Diversity benefits teams by enhancing creativity, improving decision-making, and expanding access to a broader talent pool.  76, 77. However, diversity can also have adverse effects, such as emotional conflict 78, 79, stress 80, poor work relationships, and poor overall performance 81. There is a lack of consensus on which aspects of group diversity are likely to result in positive outcomes and which aspects could produce negative outcomes 81, 82. The effects of diversity on group performance depend on the complexity of the task and various moderators, such as the density of the interaction network 83. A review underscores the need for in-depth research into surface and deep-level diversity within hybrid teams 82, exploring their impact across various outcomes and contexts over time.
Surface-level diversity